OpenStudy (anonymous):

For which operations is the set {–1, 0, 1} closed? Choose all answers that are correct. A. addition B. division C. multiplication D. subtraction

OpenStudy (anonymous):

so

OpenStudy (anonymous):

Try each operation on the elements and see what happens

OpenStudy (anonymous):

b

OpenStudy (anonymous):

going the guessing route?

OpenStudy (anonymous):

what's 1+1?

OpenStudy (agl202):

Not Division because 1/0 is undefined Not Addition because 1+1 = 2 Not Subtraction because 0-1= -1

geerky42 (geerky42):

@urbanmorgans So \(\dfrac{1}{0}\) exists in a set?

OpenStudy (agl202):

Ur last choice is?

OpenStudy (anonymous):

-1 is in the set

geerky42 (geerky42):

\(\color{blue}{\text{Originally Posted by}}\) @Agl202 Not Subtraction because 0-1= -1 \(\color{blue}{\text{End of Quote}}\) You mean -1-1, right?

OpenStudy (anonymous):

-1 - 1 is not

OpenStudy (agl202):

Ya, sorry my bad

OpenStudy (anonymous):

\[ 1 - (-1) = 2 \]

OpenStudy (anonymous):

this person doesn't seem to want to do the work to learn, though

OpenStudy (anonymous):

c